# Generative Medical Event Models Improve with Scale

### v1

@@ 2.5.1 Hospital length of stay and 30-day readmission
Accurate hospital length of stay (LOS) prediction helps health systems manage beds and plan patient care to limit discharge delays [31]. We evaluated CoMET’s ability to predict LOS on 10,000 randomly selected hospital admissions (see Table 21 for evaluation set statistics). All models received the patient’s history through the admission encounter’s header (i.e., encounter type, department specialty, and any chief complaints), in addition to the documented primary diagnosis.
Another metric for assessing healthcare utilization is the 30-day hospital readmission risk, an important and well-studied operational consideration for care transitions, discharge planning, and outpatient follow-up [32]. We randomly selected 10,000 patients being discharged from the hospital and had models predict the probability that a new hospitalization would occur within 30 days (patients who in real life were readmitted within a day were excluded from this analysis because they may often reflect hospital transfers or clerically erroneous discharges [33]). CoMET models saw increases in AUCROC on this task, with CoMET-L and CoMET-M demonstrating higher scores than task-specific supervised models (see Figure 14 and Appendix B).
Figure 12: One-year encounter frequency forecasting. CoMET models were compared with the best-performing supervised task-specific regression model for predicting the number of inpatient, emergency, and outpatient encounters that will occur within a year’s time for 18,400 patients. Mean absolute error (MAE) was used to measure the error in predicting these encounter counts; smaller the better.
Figure 13: Hospital length of stay prediction. Each of the CoMET models and the best-performing task-specific model evaluated on LOS prediction on a set of 10,000 randomly selected samples. Mean absolute error (MAE) of the length of stay in days was used to compare model performance; smaller the better.
Figure 14: 30-day readmission prediction. ROC curves for predicting hospital readmission within 30 days of discharge across 10,000 hospital encounters. CoMET-L achieved an AUCROC of 0.770, with the best-performing supervised task-specific model achieving 0.717.
@@ 2.6 Training medical event models follows scaling laws
Before training CoMET-S, CoMET-M, and CoMET-L, we first trained many smaller models to derive scaling laws [34, 35] to predict the optimal model size and number of training tokens for a given compute budget, measured in tera floating point operations. This step was important not only for understanding the best parameters for training models, but also to understand how to optimally scale medical event foundation models on a sufficiently large dataset. Building on recent work demonstrating power-law scaling for generative medical event prediction on the MIMIC-IV dataset [17], we applied the same approach to Cosmos by training a sweep of 10 model sizes ranging from two million to one billion parameters on our dataset of over 136 billion training tokens.
As in Hoffmann et al. [34], we ran a grid search over varying amounts of training TFLOPs to find the optimal scaling of model size and training tokens for the CoMET medical event dataset. After performing smaller training runs at four fixed compute budgets, we fit parabolas to the isoFLOP curves (Figure 15).
Figure 15: Optimal training of CoMET models follows power laws. The minimum loss achieved by training runs with fixed compute and varied number of model parameters (left). A log-scale parabola is fit to each isoFLOP curve, and the minimum point of each is marked by a diamond. Middle and right show how the isoFLOP parabolas’ minima $N_{\text{opt}}$ and $D_{\text{opt}}$, respectively, vary with the isoFLOP compute on a log-log scale with power-law fits.
@@
with fit parameters $A$, $B$, $\alpha$, and $\beta$, we obtained best-fit parameters of $\alpha=0.520$ and $\beta=0.512$, closely mirroring coefficients reported for natural‑language corpora ($\alpha=0.49$ and $\beta=0.51$[34]. The near-equality of $\alpha$ and $\beta$ implies that, on the compute-optimal frontier, model size and training-token count should be scaled proportionally as total compute increases. We used these best-fit lines to derive the training parameters to train our 3 compute-optimal CoMET models, listed in Table 1.
@@ 2.7 Scaling medical event model performance
Our scaling analysis demonstrated that increasing parameters and training tokens predictably decreases model loss. This raises the question of how minimized loss corresponds with downstream, clinically relevant evaluations. To measure this relationship, we evaluated the CoMET models and several training checkpoints at different train loss on the single-encounter generation, T2DM-specific outcomes, and 30-day readmissions tasks.
@@
Figure 16: Downstream performance improves as train loss decreases. We evaluated each of the CoMET models, as well as earlier checkpoints from the CoMET-S and CoMET-M training runs, on a variety of our downstream evaluations. We fit a sigmoid curve to all points except for CoMET-L to assess the sigmoid curve’s predictive utility. We evaluated all of these models using a more conservative $n=20$ simulations.
@@ 2.7.2 Scaling test-time compute
At inference time, one of the uniquely relevant parameters for CoMET is the number of generated patient timelines, $n$. Because probabilities are calculated as an aggregation over $n$ generations (Section 5.5), increasing $n$ scales inference costs linearly while reducing the Monte Carlo variance of downstream predictions and increasing the resolution of output probabilities. Unlike in language models, where test-time compute navigates a fully expressive space, in our setting, increasing $n$ also expands the method’s expressiveness, raising the performance ceiling by reducing the quantization of predictions. To investigate this tradeoff, we varied $n$ for three representative clinical tasks: hyperlipidemia-specific outcomes (an evaluation with relatively high-prevalence positive outcomes), incident disease prediction (an evaluation with the lowest-prevalence positive outcomes), and single-encounter emergency generations (which is highly multi-target and uses PR-AUC).
Figure 17: Effect of test-time compute on performance. For CoMET-L we evaluated model performance against the number of simulations generated, focusing on single-encounter generations for emergency visits (left), HLD-specific outcomes (middle), and incident disease tasks (right). Each line represents a specific task (e.g., one-year ASCVD risk, two-year COPD, etc.) from the titled category. For readability the legend is not included.
@@ 3.1 Key findings
Here we demonstrate that large‐scale, medical event foundation models can learn from longitudinal patient records to produce realistic and useful clinical predictions. Trained on a subset of the 16.3B encounters across 300M patient records as of August 2025 from Cosmos, CoMET is a family of decoder-only transformer models up to 1B parameters in size that probabilistically generate medical event sequences which not only preserve event- and encounter-level realism but also demonstrate a broad range of short- and long-term predictive capabilities.
@@
Previous medical event foundation models have been constrained by training on smaller datasets or a limited number of data types [37]. CoMET models were trained directly on sequences of structured medical events spanning core clinical data types (i.e., demographics, encounters, diagnoses, chief complaints, labs, medications, procedures, and time). Rather than relying on natural language inputs and outputs, CoMET directly receives discrete medical events from a patient’s medical record as input, learning event-level structure, long-range temporal dependencies, and the underlying probability distribution of medical events. As of this writing, CoMET represents the largest medical-event foundation model by number of medical events used for training.
@@
Evaluation of medical event foundation models has typically been performed on relatively few evaluation tasks [37]. By assessing CoMET’s performance across a broad and diverse suite of clinical tasks spanning single-encounter generation, disease-specific outcome prediction, acute-on-chronic risk estimation, incident disease detection, differential diagnosis, and operational forecasting, we explored its potential as a flexible and generalizable tool for a range of clinical contexts. CoMET-L, our largest model, matched or outperformed supervised task-specific models on most of these tasks. Performance gains were evident across multiple clinical domains, supporting CoMET’s generalizability and utility for diverse downstream applications.
@@
We also conducted the largest scaling law analysis for medical event data to date—more than $300\times$ larger than the largest previous work [17]. To investigate medical event scaling laws, we employ the same methodology from [34], exploring isoFLOPs across four fixed compute points. Our isoFLOPs analysis yields similar exponents to those shown in the natural language (NL) domain ($\alpha=0.520$ and $\beta=0.512$), suggesting we should scale training tokens and model parameters proportionally to train compute-optimal models [34]. In contrast to work in the NL domain, we observe an optimal token-to-parameter ratio of  1,000:1, closely matching results from prior work on medical event scaling laws [17].
@@
Lastly, we found that loss minimization predictably translates into better downstream performance on our suite of evaluations. Nearly every clinical task we evaluated exhibited sigmoidal improvements in their performance metrics as we minimized training loss. These trends were consistent across task families and clinical domains, and most tasks had not yet plateaued, indicating substantial remaining headroom. This result highlights pretraining loss as a useful proxy for downstream capability, and suggests that continuing to scale will yield better downstream utility. In addition to training-time scaling, we also observe benefits from increased inference-time compute: generating more patient simulations per prediction consistently improves downstream performance metrics across many clinical tasks.
@@ 3.2 Limitations
Even with these strengths, CoMET has clear limitations.
Real-world data. CoMET is trained on real-world healthcare data in Cosmos, and it is subject to imprecision and errors in documentation and clinical practice [38]. Several automated data quality control processes in Cosmos aim to improve data completeness and identify irregularities, and these periodic analyses are fed back to the contributing organizations [6]. Cosmos aggregates data across 310 health systems with extensive linking and deduplication efforts, but individuals may receive care at other health systems that is not recorded in Cosmos. This leads to data gaps before, within, and after observed medical event sequences. Certain prediction targets may be missing or occur earlier than the first record in Cosmos. We mitigate this by training and evaluating on patients who meet predefined criteria for observable healthcare utilization (see Section 5.1) and by the use of time tokens, which allow the model to read and generate patterns of absent data.
Tokenization. While the tokenizer for CoMET can be used to encode a wide range of clinical data, the discretization of continuous values like laboratory results and time masks smaller changes in these values that may be important for some prediction tasks. Future work can investigate tradeoffs in performance and model complexity between a pure categorical approach and a hybrid approach that handles numeric values separately [39, 40, 41, 42, 43]. In addition, we did not conduct an evaluation for training a tokenizer or optimizing vocabulary size, presenting a rich avenue for future work. Tokenizers trade vocabulary size for sequence length which can impact performance in non-trivial ways with model scale [44]. As shown in Figure 18, many patients’ entire tokenized history exceeded the model’s current context window of 8,192, presenting an opportunity to improve performance with longer context models. Finally, while CoMET tokenized and was trained on a core set of structured clinical data types Section 5.2, many other data types remain to be added.
Evaluation. Disease phenotype definitions were based on International Classification of Diseases, 10th Revision, Clinical Modification (ICD-10-CM) codes, categorical encounter types, demographics, event timings, and other similar features. While these features form reasonable phenotypes, they do not always correspond directly to how events transpired in the real world. This presents an opportunity for future work involving prospective validation of CoMET’s outputs. Finally, evaluations focused on aggregate performance across individuals in the test set and did not investigate performance and calibration within specific subpopulations, which will be an additional area for future work.
@@ 3.3 Future work
The above constraints outline the frontier of CoMET’s capabilities. However, each constraint naturally suggests consequent experiments and improvements.
New event types. First, including additional structured data from Cosmos, such as genomic variants, social drivers of health, and cancer staging data, would give CoMET a more robust representation of patient health timelines. The addition of specialty-focused data types would further improve CoMET’s ability to both contextualize and make certain predictions for a wide range of medical specialties, professions, and use cases. Pediatric data, including mother-baby links, could enable CoMET to capture developmental physiology, age-specific drug dosing, vaccine schedules, and congenital condition timelines. Lastly, multimodal data such as waveform vitals, images, and clinical notes would close information gaps that structured medical event data cannot capture.
Counterfactual reasoning. Future work will incorporate decision-effect estimation objectives and reinforce counterfactual consistency; explicit what-if analyses (e.g., “What if metformin were initiated today?”) will allow us to move beyond risk analysis and towards actionable risk mitigation.
Time to event analysis. Our evaluation tasks focused primarily on binary predictive outcomes, and censoring was not explicitly accounted for in our analysis of real and synthetic patient timelines. This does not take advantage of the full time-to-event predictive capacity that CoMET has to offer.
Fine-tuning. While CoMET achieves strong zero-shot performance, it can also be fine-tuned to improve task-specific performance and extend its capabilities to out of vocabulary tasks. Fine tuning offers an alternative to increasing test-time compute, eliminating sampling variance and enabling multiple precise calibrated predictions to be made with a single compute-efficient forward pass.
Prospective studies. Prospective studies of CoMET could help assess how well CoMET outputs align with clinical practice, which would inform downstream applications and model retraining schedules.
Human factors research and governance. CoMET models are generalizable to many clinical tasks, and deeply understanding best practices for introducing and scaling downstream applications within clinical workflows is key to interpretability, transparency, and usability. As with other healthcare technology, strong evaluation frameworks and thoughtful governance strategies are important to make CoMET applications sustainable, responsible, and fair.
Together, these findings position CoMET as a general‐purpose engine for real-world evidence: it can screen populations for incident disease risk, forecast individual health timelines, surface differential diagnoses, and anticipate health system interactions—all from a single set of autoregressive generations.
@@ 4 Related Work
In this section, we highlight prior work that informs and contextualizes our study, focusing on three key areas: (1) large-scale, real-world medical event databases such as Epic Cosmos, (2) scaling laws for large language models, and (3) medical foundation models trained on structured and unstructured health data. This is not intended as an exhaustive survey of all related literature; rather, we summarize representative work that captures the state of the art, key challenges, and motivations for our approach.

### v3

@@ 2.5.1 Hospital length of stay and 30-day readmission
Accurate hospital length of stay (LOS) prediction helps health systems manage beds and plan patient care to limit discharge delays [31]. We evaluated Curiosity’s ability to predict LOS on 10,000 randomly selected hospital admissions (see Table 21 for evaluation set statistics). All models received the patient’s history through the admission encounter’s header (i.e., encounter type, department specialty, and any chief complaints), in addition to the documented primary diagnosis.
Another metric for assessing healthcare utilization is the 30-day hospital readmission risk, an important and well-studied operational consideration for care transitions, discharge planning, and outpatient follow-up [32]. We randomly selected 10,000 patients being discharged from the hospital and had models predict the probability that a new hospitalization would occur within 30 days (patients who were readmitted within one day were excluded from this analysis because they may often reflect hospital transfers or clerically erroneous discharges [33]). Curiosity models saw increases in AUCROC on this task, with Curiosity-L and Curiosity-M demonstrating higher scores than task-specific supervised models (see Figure 14 and Appendix B).
Figure 12: One-year encounter frequency forecasting. Curiosity models were compared with the best-performing supervised task-specific regression model for predicting the number of inpatient, emergency, and outpatient encounters that will occur within a year’s time for 18,400 patients. Mean absolute error (MAE) was used to measure the error in predicting these encounter counts; smaller is better.
Figure 13: Hospital length of stay prediction. Each of the Curiosity models and the best-performing task-specific model evaluated on LOS prediction on a set of 10,000 randomly selected samples. Mean absolute error (MAE) of the length of stay in days was used to compare model performance; smaller is better.
Figure 14: 30-day readmission prediction. ROC curves for predicting hospital readmission within 30 days of discharge across 10,000 hospital encounters. Curiosity-L achieved an AUCROC of 0.770, with the best-performing supervised task-specific model achieving 0.717.
@@ 2.6 Training medical event models follows scaling laws
Before training Curiosity-S, Curiosity-M, and Curiosity-L, we first trained many smaller models to derive scaling laws [34, 35] to predict the optimal model size and number of training tokens for a given compute budget, measured in tera floating point operations. This step was important not only for understanding the best parameters for training models, but also to understand how to optimally scale medical event foundation models on a sufficiently large dataset. Building on recent work demonstrating power-law scaling for generative medical event prediction on the MIMIC-IV dataset [17], we applied the same approach to Cosmos by training a sweep of 10 model sizes ranging from two million to one billion parameters on our dataset of over 136 billion training tokens.
As in Hoffmann et al. [34], we ran a grid search over varying amounts of training TFLOPs to find the optimal scaling of model size and training tokens for the Curiosity medical event dataset. After performing smaller training runs at four fixed compute budgets, we fit parabolas to the isoFLOP curves (Figure 15).
Figure 15: Optimal training of Curiosity models follows power laws. The minimum loss achieved by training runs with fixed compute and varied number of model parameters (left). A log-scale parabola is fit to each isoFLOP curve, and the minimum point of each is marked by a diamond. Middle and right show how the isoFLOP parabolas’ minima $N_{\text{opt}}$ and $D_{\text{opt}}$, respectively, vary with the isoFLOP compute on a log-log scale with power-law fits.
@@
with fit parameters $A$, $B$, $\alpha$, and $\beta$, we obtained best-fit parameters of $\alpha=0.520$ and $\beta=0.512$, closely mirroring coefficients reported for natural‑language corpora ($\alpha=0.49$ and $\beta=0.51$[34]. The near-equality of $\alpha$ and $\beta$ implies that, on the compute-optimal frontier, model size and training-token count should be scaled proportionally as total compute increases. We used these best-fit lines to derive the training parameters to train our 3 compute-optimal Curiosity models, listed in Table 1.
@@ 2.7 Scaling medical event model performance
Our scaling analysis demonstrated that increasing parameters and training tokens predictably decreases model loss. This raises the question of how minimized loss corresponds with downstream, clinically relevant evaluations. To measure this relationship, we evaluated the Curiosity models and several training checkpoints at different train loss on the single-encounter generation, T2DM-specific outcomes, and 30-day readmissions tasks.
@@
Figure 16: Downstream performance improves as train loss decreases. We evaluated each of the Curiosity models, as well as earlier checkpoints from the Curiosity-S and Curiosity-M training runs, on a variety of our downstream evaluations. We fit a sigmoid curve to all points except those from Curiosity-L to assess the sigmoid curve’s predictive utility. We evaluated all of these models using a more conservative $n=20$ simulations.
@@ 2.7.2 Scaling test-time compute
At inference time, one of the uniquely relevant parameters for Curiosity is the number of generated patient timelines, $n$. Because probabilities are calculated as an aggregation over $n$ generations (Section 5.5), increasing $n$ scales inference costs linearly while reducing the Monte Carlo variance of downstream predictions and increasing the resolution of output probabilities. Unlike in language models, where test-time compute navigates a fully expressive space, in our setting, increasing $n$ also expands the method’s expressiveness, raising the performance ceiling by reducing the quantization of predictions. To investigate this tradeoff, we varied $n$ for three representative clinical tasks: hyperlipidemia-specific outcomes (an evaluation with relatively high-prevalence positive outcomes), incident disease prediction (an evaluation with the lowest-prevalence positive outcomes), and single-encounter emergency generations (which is highly multi-target and uses PR-AUC).
Figure 17: Effect of test-time compute on performance. For Curiosity-L we evaluated model performance against the number of simulations generated, focusing on single-encounter generations for emergency visits (left), HLD-specific outcomes (middle), and incident disease tasks (right). Each line represents a specific task (e.g., one-year ASCVD risk, two-year COPD, etc.) from the titled category. For readability the legend is not included.
@@ 3.1 Key findings
Here we demonstrate that large‐scale, medical event foundation models can learn from longitudinal patient records to produce realistic and useful clinical predictions. Trained on a subset of the 16.3B encounters across 300M patient records as of August 2025 from Cosmos, Curiosity is a family of decoder-only transformer models up to 1B parameters in size that probabilistically generate medical event sequences which not only preserve event- and encounter-level realism but also demonstrate a broad range of short- and long-term predictive capabilities.
@@
Previous medical event foundation models have been constrained by training on smaller datasets or a limited number of data types [37]. Curiosity models were trained directly on sequences of structured medical events spanning core clinical data types (i.e., demographics, encounters, diagnoses, chief complaints, labs, medications, procedures, and time). Rather than relying on natural language inputs and outputs, Curiosity directly receives discrete medical events from a patient’s medical record as input, learning event-level structure, long-range temporal dependencies, and the underlying probability distribution of medical events. As of this writing, Curiosity represents the largest medical-event foundation model by number of medical events used for training.
@@
Evaluation of medical event foundation models has typically been performed on relatively few tasks [37]. By assessing Curiosity’s performance across a broad and diverse suite of clinical tasks spanning single-encounter generation, disease-specific outcome prediction, acute-on-chronic risk estimation, incident disease detection, differential diagnosis, and operational forecasting, we explored its potential as a flexible and generalizable tool for a range of clinical contexts. Curiosity-L, our largest model, matched or outperformed supervised task-specific models on most of these tasks. Performance gains were evident across multiple clinical domains, supporting Curiosity’s generalizability and utility for diverse downstream applications.
@@
We also conducted the largest scaling law analysis for medical event data to date—more than $300\times$ larger than the largest previous work [17]. To investigate medical event scaling laws, we employed the same methodology from [34], exploring isoFLOPs across four fixed compute points. Our isoFLOPs analysis yielded similar exponents to those shown in the natural language (NL) domain ($\alpha=0.520$ and $\beta=0.512$), suggesting we should scale training tokens and model parameters proportionally to train compute-optimal models [34]. In contrast to work in the NL domain, we observed an optimal token-to-parameter ratio of  1,000:1, closely matching results from prior work on medical event scaling laws [17].
@@
Lastly, we found that loss minimization predictably translates into better downstream performance on our suite of evaluations. Nearly every clinical task we evaluated exhibited sigmoidal improvements in its performance metrics as we minimized training loss. These trends were consistent across task families and clinical domains, and most tasks had not yet plateaued, indicating substantial remaining room for improvement. This result highlights pretraining loss as a useful proxy for downstream capability, and suggests that continuing to scale will yield better downstream utility. In addition to training-time scaling, we also observe benefits from increased inference-time compute: generating more patient simulations per prediction consistently improves downstream performance metrics across many clinical tasks.
@@ 3.2 Limitations
Even with these strengths, Curiosity has clear limitations.
Real-world data. Curiosity is trained on real-world healthcare data in Cosmos, and it is subject to imprecision and errors in documentation and clinical practice [38]. Several automated data quality control processes in Cosmos aim to improve data completeness and identify irregularities, and these periodic analyses are fed back to the contributing organizations [6]. Cosmos aggregates data across 310 health systems with extensive linking and deduplication efforts, but individuals may receive care at other health systems that is not recorded in Cosmos. This leads to data gaps before, within, and after observed medical event sequences. Certain prediction targets may be missing or occur earlier than the first record in Cosmos. We mitigate this by training and evaluating on patients who meet predefined criteria for observable healthcare utilization (see Section 5.1) and by using time tokens, which allow the model to read and generate patterns of absent data.
Tokenization. While the tokenizer for Curiosity can be used to encode a wide range of clinical data, the discretization of continuous values like laboratory results and time masks smaller changes in these values that may be important for some prediction tasks. Future work can investigate trade-offs in performance and model complexity between a pure categorical approach and a hybrid approach that handles numeric values separately [39, 40, 41, 42, 43]. In addition, we did not conduct an evaluation for training a tokenizer or optimizing vocabulary size, presenting a rich avenue for future work. Tokenizers trade vocabulary size for sequence length, which can impact performance in nontrivial ways with model scale [44]. As shown in Figure 18, many patients’ entire tokenized history exceeded the model’s current context window of 8,192, presenting an opportunity to improve performance with longer context models. Finally, while Curiosity tokenized and was trained on a core set of structured clinical data types Section 5.2, many other data types remain to be added.
Evaluation. Disease phenotype definitions were based on International Classification of Diseases, 10th Revision, Clinical Modification (ICD-10-CM) codes, categorical encounter types, demographics, event timings, and other similar features. While these features form reasonable phenotypes, they do not always correspond directly to how events transpired in the real world. This presents an opportunity for future work involving prospective validation of Curiosity’s outputs. Finally, evaluations focused on aggregate performance across individuals in the test set and did not investigate performance and calibration within specific subpopulations, which will be an additional area for future work.
@@ 3.3 Future work
The above constraints outline the frontier of Curiosity’s capabilities. However, each constraint naturally suggests consequent experiments and improvements.
New event types. First, including additional structured data from Cosmos, such as genomic variants, social drivers of health, and cancer staging data, would give Curiosity a more robust representation of patient health timelines. The addition of specialty-focused data types would further improve Curiosity’s ability to both contextualize and make certain predictions for a wide range of medical specialties, professions, and use cases. Pediatric data, including mother-baby links, could enable Curiosity to capture developmental physiology, age-specific drug dosing, vaccine schedules, and congenital condition timelines. Lastly, multimodal data such as waveform vitals, images, and clinical notes would close information gaps that structured medical event data cannot capture.
Counterfactual reasoning. Future work will incorporate decision-effect estimation objectives and reinforce counterfactual consistency; explicit what-if analyses (e.g., “What if metformin were initiated today?”) will allow us to move beyond risk analysis and toward actionable risk mitigation.
Time-to-event analysis. Our evaluation tasks focused primarily on binary predictive outcomes, and censoring was not explicitly accounted for in our analysis of real and synthetic patient timelines. This does not take advantage of the full time-to-event predictive capacity that Curiosity has to offer.
Fine-tuning. While Curiosity achieves strong zero-shot performance, it can also be fine-tuned to improve task-specific performance and extend its capabilities to out of vocabulary tasks. Fine-tuning offers an alternative to increasing test-time compute, eliminating sampling variance and enabling multiple precise calibrated predictions to be made with a single compute-efficient forward pass.
Prospective studies. Prospective studies could help assess how well Curiosity outputs align with clinical practice, which would inform downstream applications and model retraining schedules.
Human factors research and governance. Curiosity models are generalizable to many clinical tasks, and deeply understanding best practices for introducing and scaling downstream applications within clinical workflows is key to interpretability, transparency, and usability. As with other healthcare technology, strong evaluation frameworks and thoughtful governance strategies are important to make Curiosity applications sustainable, responsible, and fair.
Together, these findings position Curiosity as a general‐purpose engine for real-world evidence: it can screen populations for incident disease risk, forecast individual health timelines, surface differential diagnoses, and anticipate health system interactions—all from a single set of autoregressive generations.
@@ 4 Related Work
In this section, we highlight prior work that informs and contextualizes our study, focusing on three key areas: (1) large-scale, real-world medical event databases such as Epic Cosmos, (2) scaling laws for large language models, and (3) medical foundation models trained on structured and unstructured health data. This is not an exhaustive survey of related literature; rather, we summarize representative work that captures the state of the art, key challenges, and motivations for our approach.